OpenStudy (anonymous):

Claire lives 6/8 of a mile from the store and 4/5 of this distance from her best friend. How far does Claire live from her best friend?

OpenStudy (johnweldon1993):

well....she lives 6/8 of a mile from the store....and 4/5 of that 6/8 from her friend so 4/5 of 6/8 means 4/5 * 6/8 6*4 = 24 8*5 = 40 24/40 ------ 8 / 8 3 / 5 so she lives 3/5 of a mile from her friend

OpenStudy (anonymous):

sure, but i need to explain my little boy here that why we multiply -- any explanation on that would be great and sincerely appreciated (he had added both the fractions!)

OpenStudy (johnweldon1993):

hmm...okay lets see Well when you add the 2 fractions.....you are adding the distances here so lets do this out that way 6 4 - + - 8 5 find the lcd which is 40 and get 30 32 -- + --- 40 40 now add to get 62 -- 40 simplify it to 31 -- 20 or as a mixed fraction 1 and 11/20 if THAT was the answer....it would be said that the distance was 1 mile and 11/20 of a mile In the question....it states that each distance....didn't even reach a mile (6/8 of a mile means less than a mile) (4/5 of that distance means an even smaller part of that so still less than a mile) so that would make no sense to add the 2 because as you can see...you wouldn't come to a logical answer so now If I asked....what is 1/2 OF 20...how would you find it? You are taking the fractional part of that number..... so 1/2 of 20......means you take 20....multiply it by 1....then divide THAT answer by 2....to get 10....you have found the fractional part (1/2) of 20. if I asked what is 2/6 of 60? again....you are finding the fractional part...of 60 for this you do the same thing...multiply the 60 by 2....then divide THAT answer by 6.....to get 20...you have found the fractional part (2/6) of 60 it works the same with fraction times a fraction here you are finding the fractional part of 6/8 so what is 4/5 of 6/8? you do the SAME thing....multiply the 2 top numbers ( 4*6) and the 2 bottom numbers (5*8) to get 24/40 now what number goes into both? 8 looks good so 24/8? = 3 40/8? = 5 so you have 3/5 you have found that 3/5 ....is the fractional part (4/5) of 6/8 I'm not sure if that helps but...